An investment project provides cash inflows of $600 per year for eight years.

  

What is the project payback period if the initial cost is $1,625? (Enter 0 if the project never pays back. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Payback period years

  

What is the project payback period if the initial cost is $3,225? (Enter 0 if the project never pays back. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Payback period years

  

What is the project payback period if the initial cost is $5,100? (Enter 0 if the project never pays back. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

    

  Payback period years
Answer #1

Payback period=initial cost/annual cash flows

1.Payback period=1625/600=2.71 years(Approx).

2.Payback period=3225/600=5.38 years(Approx).

3.Payback period=5100/600=0